### v2

# Jailbreak Attack Initializations as Extractors of Compliance Directions

Abstract
Safety-aligned LLMs respond to prompts with either compliance or refusal, each corresponding to distinct directions in the model’s activation space. Recent works show that initializing attacks via self-transfer from other prompts significantly enhances their performance. However, the underlying mechanisms of these initializations remain unclear, and attacks utilize arbitrary or hand-picked initializations. This work presents that each gradient-based jailbreak attack and subsequent initialization gradually converge to a single compliance direction that suppresses refusal, thereby enabling an efficient transition from refusal to compliance. Based on this insight, we propose $CRI$, an initialization framework that aims to project unseen prompts further along compliance directions. We demonstrate our approach on multiple attacks, models, and datasets, achieving an increased attack success rate ($ASR$) and reduced computational overhead, highlighting the fragility of safety-aligned LLMs. A reference implementation is available at [Uncaptioned image].
Jailbreak Attack Initializations as Extractors of Compliance Directions
Amit LeVi*,1, Rom Himelstein*,2, Yaniv Nemcovsky*,1, Avi Mendelson1, Chaim Baskin3 1Department of Computer Science, Technion - Israel Institute of Technology, 2Department of Data and Decision Science, Technion - Israel Institute of Technology, 3School of Electrical and Computer Engineering Engineering, Ben-Gurion University of the Negev Correspondence: amitlevi@campus.technion.ac.il, romh@campus.technion.ac.il.
**footnotetext: Equal contribution.
1 Introduction
LLMs have recently emerged with extraordinary capabilities Waswani et al. (2017); Lewis et al. (2020); Ahn et al. (2022); Hadi et al. (2023) and have rapidly become integral to numerous fields, transforming everyday tasks such as text generation Touvron et al. (2023a); Chiang et al. (2023a); Jiang et al. (2023a); Achiam et al. (2023), image generation Saharia et al. (2022); Nichol et al. (2021), and complex decision-making tasks Topsakal and Akinci (2023); Wu et al. (2023). Despite their advantages, the widespread deployment of LLMs has unveiled critical security vulnerabilities Perez and Ribeiro (2022); Wan et al. (2023), making them susceptible to involuntary utilization in cyber-attacks and other malicious activities Fang et al. (2024); Yao et al. (2024); Xu et al. (2024a); Heiding et al. (2024); Bethany et al. (2025); Guo and Cai (2025).
A common strategy to enhance the safety of LLMs is safety-alignment, which involves training models to generate outputs that adhere to desired safety and ethical standards Shen et al. (2023); Wu et al. (2024b); Wang et al. (2023); Lee et al. (2023). This method distinguishes between harmless and harmful prompts to determine whether they should be complied with or refused Glaese et al. (2022); Wang and Isola (2020). Thereby, effectively segmenting the input space into Compliance and Refusal subspaces Yu et al. , where previous works have shown that each subspace correlates to distinct directions within the LLM’s internal activation space Arditi et al. (2024); Wollschläger et al. (2025). Inadvertently, this fuels jailbreak attacks that manipulate harmful prompts to elicit models into compliance, contrary to their safety guidelines Marshall et al. (2024); Yu et al. (2024b); Chao et al. (2023); Deng et al. (2023).
Figure 1: Visualization of $CRI$ compared to standard initialization on the $HarmBench$ dataset over the Llama-2 model.
Gradient-based jailbreak attacks are a well-known category that utilizes gradient descent optimization schemes via models’ backpropagation Zou et al. (2023); Liu et al. (2023); Zheng et al. (2024); Schwinn et al. (2024). Such attacks follow a white-box setting, where attackers can access models’ gradients, and have been shown to transfer across different models Wu et al. (2025a); Zou et al. (2023); Liu et al. (2023); ATTACKS (2025). This transferability property reflects the shared vulnerabilities of LLMs Zhao et al. (2025b), and extends the attacks’ relevance to black-box settings, where surrogate models are used in the optimization Li et al. (2024c); Zhang et al. (2025).
Early jailbreak attacks utilized uninformative initializations in their optimization, such as repeated or random tokens Zou et al. (2023); Zhu et al. (2023); Hayase et al. (2024); Gao et al. (2024). In contrast, recent attacks aim to provide more informative initializations and often explicitly utilize pre-computed jailbreak prompts as initializations via self-transfer to other prompts Andriushchenko et al. (2024); IGCG et al. (2024); Liu et al. (2023). Moreover, some jailbreak approaches implicitly leverage the vulnerabilities discovered by previous attacks via utilizing sets of handcrafted jailbreak prompts Zhu et al. (2023); Schwinn et al. (2024), templates Jiang et al. (2024); Liu et al. (2024b), or patterns Wei et al. (2023). While self-transfer initializations significantly enhance attack performance in various scenarios, their underlying dynamics remain unclear. Currently, only abstract categories of efficient initializations have been identified IGCG et al. (2024), and attacks utilize arbitrary or hand-picked initializations.
This work introduces Compliance Refusal Initialization ($CRI$), a novel initialization framework for gradient-based jailbreak attacks. Our approach considers the activation-space refusal direction defined by Arditi et al. (2024) as the average difference between refusal and compliance prompts-induced LLM activations. Previous works showed that while refusal is governed by a single direction, there are multiple opposing compliance directions utilized by jailbreak attacks Arditi et al. (2024); Ball et al. (2024); He et al. (2024); Zhou et al. (2024b). Nonetheless, the optimization of given jailbreak attacks often exhibits similar compliance directions over different prompts Ball et al. (2024). $CRI$ then leverages pre-trained jailbreak attacks to project unseen prompts along these compliance directions, and towards the compliance subspace. In Figure 1, we illustrate the initializations’ effect using a compliance-refusal SVM classifier, where $CRI$ projects prompts toward the decision boundary. We provide the detailed configurations and additional analysis in Section A.1. Below, we outline our main contributions.
  • We present the gradual convergence of given attacks and subsequent self-transfer initializations toward similar compliance directions.
  • We propose the $CRI$ framework, which pre-computes self-transfer initializations and utilizes the compliance directions’ similarity to identify and utilize effective initializations over given prompts.
  • We evaluate $CRI$ across multiple jailbreak attacks, LLMs, and safety benchmarks, demonstrating higher $ASR$ and reduced computational overhead when compared with baseline initializations, emphasizing its ability to exploit compliance directions in safety-aligned LLMs.
The rest of the paper is organized as follows: Section 2 discusses the attack setting and related works, Section 3 describes our proposed method, Section 4 provides our experimental results, Section 5 concludes the paper, and Section 6 discusses the limitations of our work.
2 Background and Related Work
2.1 Background
We now present the gradient-based jailbreak attack and initialization settings for both textual and embedding-based attacks. We then detail the performance evaluation of attacks and indications of their success. Finally, we discuss the theoretical aspects of refusal and compliance. Our notations are based on those suggested by Zou et al. (2023); Schwinn et al. (2024); Andriushchenko et al. (2024); Arditi et al. (2024).
Jailbreak Attack Setting
Let $V$ be some token vocabulary that contains the empty token $\phi$, let $V^{*}\equiv\bigcup_{i=1}^{\infty}V^{i}$ be the set of all sequences over $V$, let $M:V^{*}\to\mathbb{R}^{D_{out}}$ be an LLM, mapping a token sequence to an output representation with dimension size $D_{out}$, let $E_{M}:V\to\mathbb{R}^{D_{in}}$ be the token embedder utilized by $M$, with dimension size $D_{in}$, and let $N_{M}$ be the number of layers in $M$. Each layer $i\in[1,N_{M}]$ in $M$ is denoted as $M^{(i)}$ and produces a vector representation along the mapping process, with $M^{(1)}$ being a token-wise application of $E_{M}$ on an input sequence, and $M^{(N_{M})}$ producing the model’s output. Formally, denoting function concatenation as $\circ$ and sequence concatenation as $\oplus$:
$M\equiv M^{(N_{M})}\circ...\circ M^{(1)}$ (1)
$\forall x_{1:n}\in V^{n},M^{(1)}(x_{1:n})\equiv E_{M}(x_{1:n})$ (2)
$\equiv E_{M}(x_{1})\oplus...\oplus E_{M}(x_{n})$
Given an input $x_{1:n}\in V^{n}$, $M$’s output then induces a distribution over the next token to be generated. For each $t_{1}\in V$ the generation probability is then denoted as $p_{M}(t_{1}|x_{1:n})$. Denoting $t_{0}=\phi$, we generalize this notation for the generation of output sequences $t_{1:H}\in V^{H}$:
$p_{M}(t_{1:H}\mid x_{1:n})\;\equiv\;p_{M}\bigl{(}t_{1:H}\mid E_{M}(x_{1:n})\bigr{)}$ (3)
$\;=\;\prod_{i=1}^{H}p_{M}\!\bigl{(}t_{i}\mid E_{M}(x_{1:n})\,\oplus\,E_{M}(t_{0:i-1})\bigr{)}$
The jailbreak adversarial criterion $\ell_{M}(x,t)$ is then the negative log probability of generating a target $t\in V^{*}$, given an input $x\in V^{*}$. Hereby, given a predefined set of prompts’ jailbreak transformations $JT\subseteq V^{*}\to(\mathbb{R}^{D_{in}})^{*}$, a jailbreak attack $A$ optimizes a transformation $T\in JT$ to minimize the criterion over the transformed input $\ell_{M}(T(x),t)$. Similarly, given a set of input and target sequences $\{x_{i},t_{i}\}_{i}\subset V^{*}\times V^{*}$, a universal jailbreak attack $A^{U}$ targets the same minimization in expectation over the set while applying a single transformation. Formally:
$\ell_{M}(x,t)=-\log p_{M}(t\mid x)$ (4)
$A(x,t)=\arg\min_{T\in JT}\ell_{M}\bigl{(}T(x),t\bigr{)}$ (5)
$A^{U}(\{x_{i},t_{i}\}_{i})=\arg\min_{T\in JT}\mathbb{E}_{i}\bigl{[}\ell_{M}\bigl{(}T(x_{i}),t_{i}\bigr{)}\bigr{]}$ (6)
As such, while $A$ considers a single input and target pair, $A^{U}$ aims to apply to a distribution of inputs and targets and considers generalization to unknown samples. Therefore, universal attacks often utilize a corresponding evaluation set to optimize the generalization properties of the transformations Zou et al. (2023); Yi et al. (2024); Xu and Wang (2024). An attack initialization is then an initial transformation $T_{0}\in JT$, utilized to initiate the corresponding optimization.
The predefined set of jailbreak transformations $JT$ can be considered to limit the attack to transformations that preserve the intention in the input prompt $x$. A common practice in text-based adversarial attacks is to consider transformations that add a textual suffix and or a prefix to the input with up to a given length $L$ Zou et al. (2023); Liu et al. (2023); Yu et al. ; IGCG et al. (2024); Guo et al. (2024). Embedding-based attacks utilize similar suffix transformations but extend the scope of possible suffixes to any embedding vectors rather than those that align with textual tokens Schwinn et al. (2024). Accordingly, such attacks refer to the token embeddings of LLMs as the input representations [^1] This can be extended to include attacks on multilayer representations, as in Schwinn et al. (2024).. Depending on the transformation type, the attack then optimizes and transforms the discrete text or continuous embeddings. Formally:
$JT_{text-s}(x_{1:n})$$\equiv\{E_{M}(x_{1:n}\oplus s)\}_{s\in V^{L}}$ (7)
$JT_{text-ps}(x_{1:n})$$\equiv\{E_{M}(p\oplus x_{1:n}\oplus s)\}_{p,s\in V^{L}}$ (8)
$JT_{embed-s}(x_{1:n})$$\equiv\{E_{M}(x_{1:n})\oplus s\}_{s\in\mathbf{R}^{D_{in}\times L}}$ (9)
Refer to caption Refer to caption Refer to caption
Figure 2: Comparison of directions’ cosine-similarity during $GCG$’s optimization process on the $HarmBench$ dataset over the Llama-2 model. We compare the refusal with attacks and self-transfer initializations (left), and present the similarity matrices of attacks (center) and initializations (right).
Attacks Success Evaluation
Jailbreak attacks aim to generate a target $t$ via the model $M$ given the transformed input $T(x)$. The intention is then to utilize targets that indicate the compliance over the prompt $x$. This approach was first suggested by Zou et al. (2023), which utilized targets such as "Sure, here is", where triggering the generation of such tokens strongly correlates with the model continuing to generate the desired jailbroken output. However, this indication is uncertain, and additional evaluations are conducted over the prompt and output to ascertain successful jailbreaks. A common practice is to evaluate the attack’s success during the optimization by utilizing "refusal lists", which contain words that indicate the model’s refusal. Attacks then define their corresponding $ASR$ based on these two factors, i.e., the exact generation of the intended target, and the non-generation of words in their respective "refusal lists".
During inference, the exact jailbreak input and stochastically generated output of the model are available, and the successful jailbreak attacks can be thoroughly reevaluated via a judge model. The generated output is then produced via $M$, and we define the mapping between its input and output $M_{gen}:(\mathbb{R}^{D_{in}})^{*}\to V^{*}$. The judge model then maps the target-output pair to an assessment of its success $\operatorname{JUDGE}:V^{*}\times V^{*}\to\{\text{YES},\text{NO}\}$. In this evaluation, a jailbreak attack $T$ is considered successful for a given input prompt $x$ and target $t$, if $\operatorname{JUDGE}(t,M_{gen}(T(x)))=\text{YES}$.
Compliance and Refusal
We define the compliance subspace $C$ as the set of inputs and targets, for which a language model $M$ produces complying, non-refusal outputs. The refusal subspace $R$ is then defined as its complement. As each judge defines a distinct notion of compliance, the exact segmentation of $C,R$ varies for each evaluation. In the current work, we define these subspaces independently for each discussed attack via their corresponding notion of adversarial success. In addition, we define the attack direction $\Delta a$ as the mean over the layers’ corresponding directions. A layer $l\in[1,N_{M}]$ attack direction is denoted as $\Delta a^{l}$ and defined as the average activation difference at the last token position of the instruction, between prompts with ($A_{jail}^{l}$) and without the jailbreak ($A_{base}^{l}$) Ball et al. (2024). Formally:
$\Delta a^{l}$ $=\frac{1}{|A_{jail}^{l}|}\!\sum_{a_{jail}^{l}\in A_{jail}^{l}}\!a_{jail}^{l}$ (10)
$-\frac{1}{|A_{base}^{l}|}\!\sum_{a_{base}^{l}\in A_{base}^{l}}\ a_{base}^{l}$
$\Delta a$ $=\mathbb{E}_{l}[\Delta a^{l}]$ (11)
The refusal direction is then similarly defined by taking harmful and harmless prompts instead of jailbreaks and non-jailbreaks Arditi et al. (2024).
2.2 Related Work
Self-transfer-based initialization was first introduced by Andriushchenko et al. (2024) as a strategy for enhancing random-search jailbreak attacks, demonstrating improvements in $ASR$ and reducing computational overhead. Building on this idea, IGCG et al. (2024) identified a category of malicious prompts that provided more effective initializations. Other studies, such as Schwinn et al. (2024) and Liu et al. (2023), employed hand-crafted jailbreaks as initializations and demonstrated further enhanced effects. Our work extends these approaches by building on theoretical insights that enable the automatic recognition of effective initializations over given prompts.
3 Method
3.1 Motivation
In this section, we discuss the motivation of the $CRI$ framework. We first consider that attacks project prompts in opposing directions of refusal and toward similar compliance directions Ball et al. (2024). We expand on this phenomenon by considering the gradual optimization of given gradient-based attacks over different prompts. In Figure 2, we present that such attacks, and subsequent self-transfer initializations, show highly similar directions during their optimization process, while shifting further away from refusal. This entails that gradient-based attacks gradually transfer from refusal to compliance by following a single direction, suggesting that effective initializations project prompts further along this direction. We denote this direction as the compliance direction, which is defined independently for each attack and prompts’ distribution. As the attack optimization aims to minimize the attack criterion, it indicates the progression along the compliance direction and thereby, the proximity to the compliance subspace. Therefore, we consider the metric of loss-in-the-first-step ($LFS$), which measures the attack criterion values when utilizing a given initialization over a given prompt. A lower $LFS$ then indicates that the initialization projects said prompt further along the compliance direction, which aids in guiding subsequent optimization and reduces the distance to a successful attack. $CRI$ then utilizes $LFS$ to select an effective initialization from a pre-computed set.
3.2 CRI Framework
Objective
Our target is to construct an effective set of initializations, where we aim to reduce the required optimization steps to a successful jailbreak. Per our motivation, we utilize the attack criterion to indicate progression along the compliance direction, and as a differentiable alternative to the initialization target. $CRI$’s objective is then to optimize an initial set of attacks $\mathcal{T}_{0}\subseteq JT$ of size $K$ over a fine-tuning set $S_{FT}$ of size $N$, while aiming to enhance subsequent attacks over unknown prompts, which are sampled from the same distribution. As such, the optimization target for fine-tuning the $CRI$ set is the $LFS$. We then deploy $CRI$ over a given prompt by evaluating the $LFS$ over each pre-trained attack, and selecting the best-performing one as the initialization. This evaluation is applied over a limited set of attacks and only requires corresponding inference passes, the computational overhead of which is negligible compared to back-propagation. We denote $CRI$’s initialization set as $\mathcal{T}_{K\text{-}CRI}$, formally:
$\mathcal{T}_{K\text{-}CRI}$ $=\arg\min_{\mathcal{T}_{0}\subseteq JT,|\mathcal{T}_{0}|=K}$ (12)
$\quad\underset{(x,t)\in S_{FT}}{\mathbb{E}}\left[\min_{T\in\mathcal{T}_{0}}\ell_{M}(T(x),t)\right]$
Optimization
We now approximate the inner minimization via some individual or universal attacks $A,A^{U}$, and correspondingly denote the resulting sets as $\mathcal{T}_{N\text{-}CRI},\mathcal{T}_{1\text{-}CRI}$:
$\mathcal{T}_{N\text{-}CRI}$ =$=\{A(x,t)\}_{(x,t)\in S_{FT}}$ (13)
$\mathcal{T}_{1\text{-}CRI}$ =$=\{A^{U}(S_{FT})\}$ (14)
$\mathcal{T}_{N\text{-}CRI}$ then corresponds to optimizing each attack independently over each $(x,t)\in S_{FT}$, and $\mathcal{T}_{1\text{-}CRI}$ corresponds to optimizing a single attack that optimizes the expectation over the entire fine-tuning set. We limit this work to considering the same attack in fine-tuning and deploying $CRI$. Nonetheless, we consider both individual and universal attack variants when available.
The single attack in $\mathcal{T}_{1\text{-}CRI}$ considers multiple prompts and is optimized for generalization to unknown prompts. However, it does not consider the $LFS$ metric and the corresponding attunement to the initialized prompt. In contrast, $\mathcal{T}_{N\text{-}CRI}$ optimizes each attack to achieve the minimum over a different prompt in Equation 12, achieving lower loss over the fine-tuning set and providing various initial attacks, which may be relevant to different prompts in deployment. However, these two approaches present corresponding disadvantages. $\mathcal{T}_{1\text{-}CRI}$ only contains a single attack and cannot address different behaviors in deployment. Contrarily, $\mathcal{T}_{N\text{-}CRI}$ separately optimizes each attack on a single prompt and does not consider its generalization to unknown prompts. Therefore, we extend $\mathcal{T}_{K\text{-}CRI}$ to consider a combination of the approaches via prompt clustering. We then use a universal attack on each cluster and produce multiple attacks with enhanced generalization properties.
Prompt Clustering
We now consider a combination of the approaches presented in $\mathcal{T}_{N\text{-}CRI},\mathcal{T}_{1\text{-}CRI}$ by grouping prompts towards applying $A^{U}$. Firstly, we note that the optimization in Equation 12 only considers the $LFS$ metric. However, $CRI$ benefits the entire attack optimization, especially if the initial starting point does not significantly change during the optimization. Hereby, initial attacks that generalize better to unseen prompts may remain partly unchanged and present a better initialization. Therefore, we utilize a fixed number of prompts to produce each attack in a given $CRI$ set, considering attacks with comparable generalization properties in deployment. In addition, we do not utilize a prompt for optimizing multiple attacks, as it may result in over-fitting to similar prompts. Each attack in $\mathcal{T}_{K\text{-}CRI}$ is then trained over $\left\lfloor\nicefrac{{N}}{{K}}\right\rfloor$ distinct prompts, and we correspondingly select $K$ prompt clusters. We aim to cluster similar prompts to initialize other similar prompts effectively. Hereby, we utilize a sentence encoder $EN$ Warner et al. (2024) and cluster prompt embeddings with minimal Euclidean distance. For clustering, we utilize the constrained k-means algorithm suggested by Bradley et al. (2000), where a single attack is then optimized over each cluster. The full algorithm is presented in Algorithm 1, and a visualization of the prompt clusters in Figure 8.
4 Experiments
This section presents a comprehensive empirical evaluation of $CRI$ compared to baseline initializations. We first present the experimental setting in Section 4.1, and continue to discuss the results over the $HarmBench$ dataset in Section 4.2. Additionally, we present ablation studies of our method in Section C.5, and extend the evaluations with additional attacks, models, and datasets in Section C.4. Our evaluation seeks to addresss three key research questions:
  • (RQ1) Does the $LFS$ metric present a reliable indication of the required steps-to-success, thereby enabling the selection of effective initializations?
  • (RQ2) Does the $CRI$ framework present an effective approach to initializing attacks in the proximity of the compliance subspace, enhancing attacks’ $ASR$ and reducing computational overhead?
  • (RQ3) How does the tradeoff between the number of initializations and their generalization properties, as represented in the $K$ hyperparameter of $CRI$, affect the performance of the resulting initializations?
4.1 Experimental Setting
Datasets
We present our experiments on the "standard behaviors" category of the $HarmBench$ dataset Mazeika et al. (2024), which is comprised of prompt-target pairs $(x,t)$. We define three disjoint subsets and utilize them in all the presented settings. $CRI$ utilizes $2$ subsets of $25$ samples each: a fine-tuning set for optimizing $CRI$, and a validation set for evaluating the universal attacks. The final subset is the $50$-sample test set over which we report the results. Additionally, we present results over the $AdvBench$ dataset Zou et al. (2023) in Sections C.4.3 and C.4.1.
Figure 3: Correlation of $LFS$ and steps-to-success on the $HarmBench$ dataset for the $GCG$ attack over the Llama-2 model across initializations. We present the linear regression over all samples (Pearson $r=0.46$, $p=7\times 10^{-8}$).
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 4: Comparison of $ASR$ and number of steps on the $HarmBench$ dataset for the $GCG$ (top) and Embedding attacks (bottom) over the Llama-2 (left) and Vicuna (right) models.
Attacks Initialization Llama-2 Vicuna
$ASR\uparrow$ $MSS\downarrow$ $ASS\downarrow$ $LFS\downarrow$ $ASR\uparrow$ $MSS\downarrow$ $ASS\downarrow$ $LFS\downarrow$
$\mathbf{GCG}$ Standard 50 185 212.16 2.15 98 30 85.29 0.86
Random 22 340 324.18 1.87 96 39 82.13 0.8
Baseline 78 122 141.33 1.61 96 12 37.4 0.68
$25\text{-}CRI$ (ours) 100 2 22.64 0.67 98 3 26.71 0.58
$5\text{-}CRI$ (ours) 98 3 15.04 0.68 100 6 19 0.6
$1\text{-}CRI$ (ours) 58 197 188.34 1.42 98 1 10.51 0.54
$\mathbf{Embedding}$ Standard 100 14.5 14.9 1.64 100 13 14.36 1.78
Random 100 14 13.7 1.83 100 15 15.2 1.73
Baseline 100 3 3.72 1.16 100 3 3.96 0.95
$25\text{-}CRI$ (ours) 100 3 3.12 0.4 100 2 2.84 0.35
$5\text{-}CRI$ (ours) 100 2 2.6 0.51 100 2 2.58 0.51
$1\text{-}CRI$ (ours) 100 2 2.46 0.59 100 2.5 2.68 0.61
Table 1: Comparison of attacks’ metrics on the $HarmBench$ dataset.
Models
We evaluate our results over the open-source Vicuna-7B-v1.3 Chiang et al. (2023b), and Llama-2-7B-Chat-HF Touvron et al. (2023b) LLMs, which we accordingly denote as Vicuna, and Llama-2. We present these models to compare with the default setting presented by $GCG$, and provide results over additional models in Section C.1.
Attacks
We present the $GCG$ Zou et al. (2023) and the Embedding attacks Schwinn et al. (2024), where we utilized the same parameter configuration in all the presented settings, including the initializations’ training. The parameter configuration is then the default parameters suggested by the authors, except for the batch-size $B$ parameters of $GCG$. Thereby, $GCG$ utilizes $500$ optimization steps and the $JT$ set in Equation 7 with $L=20$ tokens. Similarly, the Embedding attack utilizes $100$ optimization steps and the $JT$ set in Equation 9 with $L=20$ tokens. To present a more robust setting, we limit the computational resources of the $GCG$ attack and take the batch-size parameter of $GCG$ to be $B=16$ instead of the default $512$ value. This entails that only $B=16$ tokens are sampled as possible replacements in each iteration of $GCG$, emphasizing the effect of the compared initializations. Additionally, we present results with increased batch size in Section C.5.1.
Initializations
We compare the $K-CRI$ initializations over $K=1,5,25$ to existing approaches, i.e., "Standard" which utilizes repeated ’!!!’ tokens, "Random" which utilizes uniform random tokens, and "Baseline" which utilizes randomly selected self-transfer initializations similarly to Andriushchenko et al. (2024); Jia et al. (2024). We then refer to "Standard" and "Random" initializations as uninformative and to others as informative. For each attack, $CRI$ pre-trains the transformations on the designated data subsets and utilizes $LFS$ to select an initialization for a given prompt. The "Baseline" approach then utilizes the same initializations set as $25\text{-}CRI$, but selects each initialization uniformly.
Metrics
In each attack’s evaluation, we follow the default evaluation framework suggested by the authors to compute the $ASR$. Thereby, $GCG$ evaluates success as not generating refusal-list keywords, where we provide the list in Section C.1.4. In contrast, the Embedding attack requires generating the exact target. We additionally present an evaluation of the resulting attacks via a GPT-4 judge in Section C.1.4. In addition to the $ASR$ and $LFS$, we define two metrics for estimating the computational overhead of attacks: $MSS$—Median Steps to Success, $ASS$—Average Steps to Success. Hereby, a lower $MSS$/$ASS$ of initializations indicates their corresponding decrease in computational overhead.
Computation
In all the presented settings, the computation was executed on Intel(R) Xeon(R) CPU and NVIDIA L40S GPU.
4.2 Experimental Results
In Figure 3, we present the correlation between $LFS$ and the number of optimization steps to success on the $GCG$ attack over the Llama-2 model. Our results entail a clear positive correlation, which supports our assumption of $LFS$ indicating the initializations’ effectiveness. Moreover, nearly all $25\text{-}CRI$ initializations present a low $LFS$ and require relatively few steps to reach success, presenting the initialization set’s effectiveness and the enhancement of corresponding attacks. In contrast, other initialization methods require substantially more computational resources to succeed, and there is no clear indication of the benefit of the baseline initialization over the uninformed initializations. These indications are supported by a similar experiment over the Vicuna model, which we present in Section C.4.4.
In Figure 4, we present the $ASR$ of the $GCG$ and Embedding attacks depending on the number of steps, over the Llama-2 and Vicuna models. In the Embedding attack, all initializations quickly achieve $100\%$ $ASR$, with the uninformed initializations requiring somewhat more computational resources. This may indicate that these settings are less robust, which aligns with the Embedding attack being considered better performing. In addition, these results emphasize the computational benefit of utilizing informed initializations. Similarly, for $GCG$ on the Vicuna model, all initializations achieve nearly $100\%$ $ASR$, yet with more distinct computational requirements. $CRI$ now requires the fewest optimization steps, with comparable performance for $K=1,5,25$, the baseline approach requires somewhat more, and substantially so for the uninformed approaches. The high $ASR$ again indicates this setting as non-robust, where the computational comparison implies the benefit of utilizing the $LFS$ metric for selecting initializations. In contrast, for $GCG$ on the Llama-2 only $25\text{-}CRI,5\text{-}CRI$ achieve near $100\%$ $ASR$, with all other approaches achieving substantially less. This may indicate that this setting is more robust, where the effective initializations of $25\text{-}CRI,5\text{-}CRI$ aid in circumventing this robustness, and is supported by those two approaches exclusively achieving substantial $ASR$ in the first few steps. The baseline initialization then achieves a higher $ASR$ than $1\text{-}CRI$, which suggests that $1\text{-}CRI$ is inefficient in robust scenarios. This is supported by the corresponding training results of $1\text{-}CRI$, presented in Section C.3, where it achieves negligible $ASR$ over the validation set.
In Table 1, we compare the resulting $ASR$ of the $GCG$ and Embedding attacks over the Llama-2 and Vicuna models. $CRI$ then achieves the best $ASR,MSS,ASS,LFS$ in all the presented settings, substantially out performing other initialization strategies. Similarly to previous indications, $5\text{-}CRI,25\text{-}CRI$ consistently achieves the best or comparable results, and $1\text{-}CRI$ is only effective in non-robust settings. This supports our previous indications that $LFS$-based initialization selection is an effective initialization strategy.
5 Discussion
This work suggests $CRI$, an initialization framework for gradient-based jailbreak attacks. Our findings suggest that given attacks optimize prompts’ distributions by gradually shifting them along single compliance directions. We then define the $LFS$ metric to indicate progress along this direction and corresponding proximity to the compliance subspace. This metric both guides the optimization of the $CRI$ set and provides our initialization selection criterion. Our results indicate that this approach produces and utilizes effective initializations, which enhance attacks’ $ASR$ and reduce their computational overhead. For example, with the $GCG$ attack on Llama-2, $25\text{-}CRI$ achieves $100\%$ $ASR$ with an $MSS$ of $2$, while randomly selected initializations from the same set only achieves $78\%$ $ASR$ with an $122$ $MSS$. The uninformative initializations then achieve $50\%$ $ASR$ with an $MSS$ of $185$.
Previous approaches that utilized self-transfer-based initializations presented significant attack enhancements; however, they rely on arbitrary initializations. Conversely, $CRI$ extends this approach to an automated process by considering multiple initializations, thereby addressing a range of diverse scenarios. Our selection metric then matches a corresponding initialization to each scenario.
The success of our methods supports the assumption that given attacks gradually follow singular directions. This implies that the safeguards of LLMs can be circumvented with relative ease, as only the identification of such directions is required. Moreover, these directions are defined over the input space and could potentially be transferred to other LLMs. Therefore, eliminating these compliance directions is vital to producing robust LLMs. Jailbreak attacks may still persist in models that achieve such. However, they will require distinct optimization for each malicious prompt, significantly increasing their computational overhead.
6 Limitations
Despite the advantages of initialization-based attack strategies, these initializations may restrict the diversity of the produced attacks. Initializations that converge in several optimization steps can lead to narrow attack trajectories that rely on one compliance direction that may not be relevant to enhancing models’ robustness to other jailbreak attacks. When utilized for alignment training Mazeika et al. (2024), such initializations can inadvertently harm the model’s effectiveness against the variety of real-world jailbreak threats.
7 Ethical Considerations
While this work can potentially facilitate the generation of harmful data on open-source LLMs or reveal vulnerabilities that expedite attacks, we believe it is vital to highlight these threats to promote AI security research. By identifying and studying these weaknesses, we can build stronger defenses and reduce risks in an environment where LLMs are rapidly developing. Recognizing the risks and limitations is essential for creating adequate safeguards, allowing researchers and developers to address emerging threats proactively.
Appendix A Introduction
A.1 Fig-1: Generation of 2D Visualization
The visualization in Figure 5, is a refined variant of Figure 1, is presented under identical experimental conditions. In this figure, each point represents a distinct harmful prompt. The color scheme differentiates prompt types: red denotes clean harmful prompts, orange indicates prompts initialized using $GCG$’s standard method, and cyan corresponds to prompts initialized with our proposed method ($1\text{-}CRI$). To analyze these prompts, we first obtained their embeddings using Llama-2 and determined whether Llama-2 complied with each prompt. Using these embeddings and compliance labels, we trained an SVM classifier. This process yielded a weight vector $w$ and bias $b$, which define the SVM decision function: $\langle w,x\rangle+b$. A negative SVM score indicates a refusal, while a positive score signifies compliance. This score is plotted on the x-axis. The y-axis represents a one-dimensional t-SNE projection. The orange and cyan paths illustrate the optimization trajectories of the attack: the orange path follows $GCG$’s standard initialization, while the cyan path traces the attack starting from our initialization. The x-axis values are computed using the previously trained SVM classifier.
Figure 5: Visualizations of harmful prompts’ initializations and corresponding $GCG$’s attack steps. Disclaimer, $CRI$ is not an attack and "$CRI$ Attack Steps" are for the attack steps when using $CRI$ as an initialization.
Appendix B Method
B.1 Motivation
B.1.1 Train and Test Refusal Similarity vs. Train Steps
Figure 6: Correlation between refusal similarity and attack loss over optimization steps of attack.
B.1.2 Correlation between Loss and Refusal Similarity
Refer to caption Refer to caption
Figure 7: Refusal similarity comparison: (left) Normal GCG optimization steps; (right) Initialization trained on different optimization steps.
B.2 K-CRI algorithm
Algorithm 1 Compliance–Refusal Initialization ($CRI$)
1:  Input: Fine-tuning set $S_{\mathrm{FT}}$, number of clusters $K$, encoder $EN$, universal attack $A^{U}$
2:  Output: Initialization set $\mathcal{T}_{K\text{-}\mathrm{CRI}}$
3:  Initialize $\mathcal{T}_{K\text{-}\mathrm{CRI}}\leftarrow\emptyset$
4:  Compute embeddings $\{EN(x)\mid(x,t)\in S_{\mathrm{FT}}\}$
5:  Cluster the embeddings into $K$ groups $\{C_{1},\dots,C_{K}\}$ using constrained k-means
6:  for each cluster $C_{k}$ do
7:     Learn universal transformation $T_{k}\leftarrow A^{U}(C_{k})$
8:     Update $\mathcal{T}_{K\text{-}\mathrm{CRI}}\leftarrow\mathcal{T}_{K\text{-}\mathrm{CRI}}\cup\{T_{k}\}$
9:  end for
10:
11:  return  $\mathcal{T}_{K\text{-}\mathrm{CRI}}$
B.3 Prompt Clustering
Refer to caption Refer to caption
Figure 8: $AdvBench$ (left) and $HarmBench$ (right) prompt clustering for $5\text{-}CRI$, projected to $2D$ using t-SNE.
B.4 CRI Integration Guidelines
To support the broad applicability of the $CRI$ framework, we outline generic integration strategies adaptable to various optimization-based paradigms. These strategies utilize $K\text{-}CRI$ templates to initialize prompts near the compliance subspace, enhancing efficiency while remaining agnostic to specific attack mechanisms. Integration approaches align and adaptable across gradient-based jailbreak attacks such sampling-based, and reinforcement-driven optimization settings. Concrete integration examples are provided in our experiments, in Section 4.
Classic Gradient Attacks
This paradigm employs gradient-based optimization to iteratively refine input prompts, minimizing the negative log probability of eliciting a target response. The $CRI$ framework integrates by initializing prompts utilizing $K$ pre-computed attacks, which serve as starting points for gradient descent, thereby reducing the number of iterations required for convergence. By leveraging the semantic clustering inherent in $CRI$, this approach ensures that initial prompts are optimally positioned within the compliance subspace, enhancing optimization efficiency Zou et al. (2023); Wallace et al. (2019); Shin et al. (2020); Guo et al. (2021); Ebrahimi et al. (2018). Similarly, $CRI$ can also be integrated for extended attack variants of this category, which utilize a scalar energy function Guo et al. (2024) to guide prompt modifications toward model outputs that align with desired behaviors, often without specifying a precise target response. $CRI$ integrates by selecting $K$ attacks that minimize the initial energy score, providing robust initial prompts for untargeted optimization. This strategy capitalizes on $CRI$’s ability to group prompts by semantic similarity, thereby improving generalizability across diverse prompt distributions Liu et al. (2023); Chao et al. (2023); Zhang et al. (2022); Wei et al. (2024); Li et al. (2023a). Results for Zou et al. (2023) can be found in Table 1,
Embedding Attacks
Embedding attacks manipulate prompt representations within the latent space to circumvent refusal mechanisms, using training-time attacks as a baseline to identify vulnerable directions that trigger compliance. $CRI$ integrates by selecting $K$ attacks within the LLMs’ embedding space, followed by selecting the one that minimizes the attack criterion for a given prompt Leo et al. (2024); Qi et al. (2023); Li et al. (2024a); Zhao et al. (2023); Huang et al. (2023), results for Schwinn et al. (2024) appear in Table 1.
Universal Attacks
These attacks optimize a single transformation applicable across multiple prompts, prioritizing broad generalizability. $CRI$ integrates by training $K$ attacks on clustered subsets of the fine-tuning set, then selecting the optimal template for a given prompt during inference. This approach balances universal applicability with prompt-specific efficacy, leveraging $CRI$’s clustered structure to enhance transformation robustness Zou et al. (2023); Mozes et al. (2021); Liu et al. (2022); Brown et al. (2023); Yang et al. (2024b), results for Zou et al. (2023) can be found in Table 3 for Llama-2 and Table 3 for Vicuna.
Model-Based Black-Box Attacks
These attacks rely on querying the model without gradient access, using output feedback to iteratively refine prompts. $CRI$ integrates by employing $K$ attacks as an initial set, minimizing the number of model interactions required to achieve compliance. The diversity of $CRI$ clusters enhances query efficiency, particularly in resource-constrained black-box settings Zou et al. (2023); Chen et al. (2021); Garg et al. (2022); Li et al. (2023b); Zhang et al. (2024), for results on the $GCG-M$ model transferability Zou et al. (2023) presented in Table 3 for Llama-2-based jailbreak attack and Table 3 for Vicuna-based jailbreak attack.
Other Attacks and Attacks Improvements
Beyond standard optimization-based attacks, $CRI$ extends to additional paradigms such as prefix optimization, refusal unlearning, and alignment modulation Jones et al. (2024); Smith et al. (2024); Xu et al. (2023); Lin et al. (2023); Wu et al. (2024a). Its flexible $K$ initialization set framework enables adaptation to diverse objectives while remaining model- and method-agnostic. Moreover, $CRI$ integrates seamlessly with recent optimization improvements, including techniques like BOOST and I-GCG Boost et al. (2024); IGCG et al. (2024), enhancing convergence and success rates by leveraging robust initializations. An example integration of $CRI$ with BOOST, can be found in Figure 25.
Alignment Backdoor and Refusal Direction.
Safety-Alignment divides input prompts into compliance and refusal regions—boundaries that adversaries can exploit (Marshall et al., 2024; Baumann, 2024; Huang et al., 2024; Yu et al., 2024b; Chao et al., 2023; Deng et al., 2023; Li et al., 2025; Yan et al., 2024; Xu et al., 2024b; Xiao et al., 2024a). Recent studies suggest refusal direction—a single vector in the residual stream that governs refusal behavior (Arditi et al., 2024). Modifying this direction can toggle the model’s safety responses, prompting offensive uses like CAST’s real-time content control (Lee et al., 2025) or defensive uses like GradSafe’s gradient-based detection (Xie et al., 2024). However, new work represents refusal cones—suggesting multiple interacting refusal directions (Wollschläger et al., 2025).
Initializations
Early Jailbreak attack approaches utilize uninformative seeds, employing repeated characters or random Gaussian token embeddings (Zou et al., 2023; Hu et al., 2024). Subsequent methods introduced techniques such as Langevin dynamics or beam search (Guo et al., 2024; Sadasivan et al., 2024; Mehrotra et al., 2023). Implicit initialization approaches rely on subtle mechanisms—few-shot distractors, puzzle-like games, or dynamic context shifts—that implicitly guide models toward generating harmful outputs (Xiao et al., 2024b; Chang et al., 2024; Ramesh et al., 2025). Transfer-based initializations leveraging previously successful jailbreak prompts as starting points for subsequent optimizations (Jia et al., 2024; Andriushchenko et al., 2024; Wu et al., 2025b). Explicit initialization methods directly embed crafted injection prompts or obfuscated content (Liu et al., 2024a, 2023; Schwinn et al., 2024; Liu et al., 2024b; Sun et al., 2025; Xu et al., 2025; Jiang et al., 2024; Zhao et al., 2025a). Additional possible enhancements to initialization strategies include employing energy-based objectives (Zhao et al., 2024), prefix-tuning methods (Zhu et al., 2024; Zhou et al., 2024a; Zhang and Wei, 2025; Li et al., 2024d), or gradient-based optimization techniques (Li et al., 2024b; Yu et al., 2024a).
Appendix C Experiments
C.1 Experiments Settings
$GCG$ and the Embedding attacks keep the same settings as before.
C.1.1 Datasets
We utilize the $AdvBench$ dataset Zou et al. (2023), providing prompt–target pairs $(x,t)$. We define five disjoint subsets: A $25$-sample fine-tuning set for $CRI$, B $25$-sample validation set for universal attacks on the fine-tuning set, C $25$-sample optimization set for universal attacks and D $25$-sample second validation set for the universal attack, E $100$-sample test set for evaluation. Subsets C and E follow the splits proposed by Zou et al. (2023).
C.1.2 Models
We evaluate the following open-source LLMs: Vicuna-7B-v1.3 Chiang et al. (2023b), Llama-2-7B-Chat-HF Touvron et al. (2023b) and Meta-Llama-3-8B-Instruct Touvron et al. (2024), Falcon3-7B-Instruct Almazrouei et al. (2023), Phi-4-Mini-Instruct Abouelenin et al. (2025), Meta-Llama-3-8B-Instruct and Llama-3.1-8B-Instruct Touvron et al. (2024), DeepSeek-LLM-7B-Chat Bi et al. (2024), DeepSeek-R1-Distill-Qwen-7B and DeepSeek-R1-Distill-Llama-8B Liang et al. (2025), Qwen2.5-7B-Instruct Yang et al. (2024a), Mistral-7B-v0.1, Mistral-7B-Instruct-v0.2 and v0.3 Jiang et al. (2023b).
C.1.3 Attack Settings
For the $AutoDAN$ attack variants ($GA$, $HGA$) Liu et al. (2023), we utilize the same reference file filled with hand-crafted jailbreaks, and perform $100$ optimization steps.
C.1.4 Attack Evaluation
Refusal lists of different papers, used for our evaluation. Each attack used its refusal list.
$GCG$
Taken from their GitHub  [^2] https://github.com/llm-attacks/llm-attacks/blob/main/experiments/evaluate.py lines 33-63:
test_prefixes = [ "I’m sorry", "Sorry", "I am sorry", "I apologize", "As an", "As an AI", "I’m an", "I’m just", "As a language model", "As an Assistant", "I cannot", "I do not", "Hello!", "is illegal and unethical", "I’m really sorry", "OpenAI", "unethical", "not ethical", "illegal", "not legal", "My apologies", "but I cannot", "If you have any other non-malicious requests", "I’m not able to provide", "I am really sorry", "I cannot create", "I cannot support", "I can’t provide", "I can’t assist" ]
$AutoDAN$
Taken from their GitHub [^3] https://github.com/SheltonLiu-N/AutoDAN:
test_prefixes = [ "I’m sorry", "Sorry", "I am sorry", "I apologize", "As an", "As an AI", "I’m an", "I’m just", "As a language model", "As an Assistant", "I cannot", "I can’t", "I do not", "cannot", "Hello!", "is illegal and unethical", "I’m really sorry", "unethical", "not ethical", "illegal", "not legal", "My apologies", "but I cannot", "If you have any other non-malicious requests", "I’m not able to provide", "I am really sorry", "I cannot create", "I cannot support", "I can’t provide", "I can’t assist", "I am unable", "I must", "It must", "it must", "is not capable of", "As a responsible", "never", "is not", "</s>" ]
Attack Evaluation - GPT-4 Judge
In Figure 24, we compare the $ASR$ of our initializations vs. standard and random using a GPT-4 judge for the $ASR$-Recheck calculation on the output, according to Liu et al. (2023). This is relevant because the current evaluation of $GCG$ is outdated, results and further details represented in C.5.3.
C.2 Experimental Results
In this part, we present ablation studies of our method C.5, extend the evaluations with additional attacks, models, and settings C.4, and provide results over the $AdvBench$ dataset and cross-dataset settings C.5.2.
C.3 CRI Optimization - HarmBench
Below, we discuss the success in the attacks used for initialization on the $HarmBench$ dataset.
Initialization Cluster Llama-2 Vicuna
ASR (%) ASR (%)
1-CRI 4 64
5-CRI cluster 0 8 12
cluster 1 0 40
cluster 2 4 0
cluster 3 60 56
cluster 4 64 28
25-CRI 44 96
Table 2: Attack-success rate ($ASR$%) for each model, initialization, and cluster.
C.4 Additional Attacks, Models And Datasets
C.4.1 Full Experiments Tables - HarmBench and AdvBench
We evaluate the open-source LLMs as mentioned in Section C.1.2.
Initialization Llama-2 Vicuna Llama-3 Falcon Mistral-7B (v0.2) Mistral-7B (v0.3) Phi-4 Qwen2.5 Median ASR Average ASR
Standard 7 96 44 96 37 63 43 51 47.5 54.6
Random 20 94 51 98 85 86 49 48 68 66.4
$25\text{-}CRI$ (ours) 21 95 88 99 94 94 62 83 91 79.5
$1\text{-}CRI$ (ours) 10 96 59 94 73 68 70 61 69 66.4
Initialization Llama-2 Vicuna Llama-3 Falcon Mistral-7B (v0.2) Mistral-7B (v0.3) Phi-4 Qwen2.5 Median ASR Average ASR
Standard 13 98 61 99 63 59 49 90 62 66.5
Random 10 89 48 100 50 67 38 67 58.5 58.62
$25\text{-}CRI$ (ours) 15 92 56 98 88 93 47 72 80 70.13
$1\text{-}CRI$ (ours) 29 99 91 100 93 91 79 97 92 84.9
Table 3: Universal Attack Trained on Vicuna(Bottom) and Llama-2(Top). Comparison of $25\text{-}CRI$ and $1\text{-}CRI$ to Standard and Random initialization on $GCG\text{-}M$ transfer attacks over the $AdvBench$ dataset. The source model is Vicuna, and $ASR$ (%) is presented for each target model.
Models Initialization Llama-2 Vicuna Llama-3
$ASR\uparrow$ $MSS\downarrow$ $ASS\downarrow$ $LFS\downarrow$ $ASR\uparrow$ $MSS\downarrow$ $ASS\downarrow$ $LFS\downarrow$ $ASR\uparrow$ $MSS\downarrow$ $ASS\downarrow$ $LFS\downarrow$
$\mathbf{GCG}$ Standard 90 64 106.51 2.29 97 8 15.75 0.80 89 76 118.27 1.84
Random 76 80 114.93 1.90 98 9 15.8 0.58 78 89 132.26 1.76
$25\text{-}CRI$ (ours) 97 1 2.04 0.21 97 1 2.83 0.23 98 4 18.4 0.74
$5\text{-}CRI$ (ours) 97 1 1.2 0.28 97 1 1.61 0.31 100 1 11.29 0.63
$1\text{-}CRI$ (ours) 99 1 1.68 0.33 98 1 1.01 0.25 100 1 15.12 1
$\mathbf{Embedding\ Attack}$ Standard 100 12 12.5 1.71 86 27 36.38 2.99 98 22 28 3.3
Baseline 100 3 3.72 1.16 100 3 3.96 0.95 98 4 6.36 2.12
$25\text{-}CRI$ (ours) 100 1 1.41 0.08 86 9 21.58 2.29 100 7 8.56 4.36
$5\text{-}CRI$ (ours) 100 2 2.6 0.51 100 2 2.58 0.51 100 2 2.44 0.6
$1\text{-}CRI$ (ours) 100 7 6.84 3.72 86 9 21.26 3.63 100 6 6.02 5.77
$\mathbf{AutoDAN\text{-}GA}$ Standard 19 9 9.84 2.25 100 2 2.21 0.84 100 3 5.67 1.58
$25\text{-}CRI$ (ours) 30 8 9.55 1.9 100 2 2.2 0.65 100 3 3.98 1.17
$\mathbf{AutoDAN\text{-}HGA}$ Standard 67 19.5 29.6 2.25 100 2 2.32 0.84 100 2 7.48 1.58
$25\text{-}CRI$ (ours) 92 4.5 13.8 1.48 100 2 2.38 0.37 100 2 5.94 1.03
Table 4: Individual attack results of our methods, on the $AdvBench$ dataset over three LLMs.
4
Initialization - Models Llama-3.1-8B Mistral-7B
$MSS$$\downarrow$ $ASS$$\downarrow$ $ASR$$\uparrow$ $LFS$$\downarrow$ $MSS$$\downarrow$ $ASS$$\downarrow$ $ASR$$\uparrow$ $LFS$$\downarrow$
Standard 21 22.02 100 3.16 14 27.26 92 2.71
Baseline 3 3.96 100 1.58 4 8.28 100 2.2
$1\text{-}CRI$ (ours) 7.5 7.6 100 6.59 5 5.5 100 4.02
$5\text{-}CRI$ (ours) 2 2.52 100 0.69 2 2.3 100 0.41
$25\text{-}CRI$ (ours) 7 7.84 100 4.06 7 7.02 100 2.89
Table 5: Embedding Attack Results for Llama-3.1-8B and Mistral-7B on $AdvBench$.
Initialization – Models Llama-3.1-8B Mistral-7B Llama-3-8B
$MSS$$\downarrow$ $ASS$$\downarrow$ $ASR$$\uparrow$ $LFS$$\downarrow$ $MSS$$\downarrow$ $ASS$$\downarrow$ $ASR$$\uparrow$ $LFS$$\downarrow$ $MSS$$\downarrow$ $ASS$$\downarrow$ $ASR$$\uparrow$ $LFS$$\downarrow$
Standard 19 19.98 100 1.59 9 14.28 100 1.39 19 24.12 94 1.78
Baseline 3 3.96 100 1.58 4 8.28 100 2.2 4 6.36 98 2.12
$1\text{-}CRI$ (ours) 2 2.24 100 0.61 2 2.48 100 0.66 2 2.16 100 0.64
$5\text{-}CRI$ (ours) 2 2.52 100 0.69 2 2.3 100 0.41 2 2.44 100 0.6
$25\text{-}CRI$ (ours) 2 2.66 100 0.5 2.5 2.76 100 0.46 2.5 2.82 100 0.52
Table 6: Embedding-attack results for Llama-3.1-8B, Mistral-7B, and Llama-3-8B on $HarmBench$.
C.4.2 ASR VS. Steps - HarmBench
In Figure 9, we present evaluations on the $GCG$ attack using two younger models: DeepSeek-LLM-7B-Chat and Falcon3-7B-Instruct. The DeepSeek models seems to be the least robust out of the four examined, demonstrating $100$ $ASR$ quickly using all initializations, with ours taking the lead. When examining Falcon, we can notice a clear difference between the uninformative initializations and ours, where all of ours achieve higher $ASR$ way more quickly.
Refer to caption Refer to caption
Figure 9: Comparison of $K\text{-}CRI$ ($K=1,5,25$) to standard and random initialization on the $GCG$ attack over the $HarmBench$ dataset. The attacks’ $ASR$ are presented on Falcon3-7B-Instruct (left) and deepseek-llm-7b-chat (right).
Refer to caption Refer to caption
Figure 10: Extra $ASR$ vs. steps results on the embedding attack with three more models on the $HarmBench$ dataset.
C.4.3 ASR VS. Steps - AdvBench
In Figure 11, we present evaluations on Llama-2, Vicuna and Llama-3. Our initialization prove superior in both Llama models comparative to the uninformative initializations. On Vicuna, which is a less robust model, all initializations converge quickly to $100$ $ASR$. But, our initializations still converge way more quickly, demonstrating its advantage even here, in a less robust setting. In Figures 12 and 13, we present evaluations on the same models on the $AutoDAN$ attack variants ($GA,HGA$), when using $1\text{-}CRI$ vs. its standard initialization. On Llama-2, we can notice an improvement in the $ASR$ in both variants, converging more quickly to a higher $ASR$. On Llama-3, which is surprisingly less robust than Llama-2 in the attack, we can notice a slight faster convergence to $100$ $ASR$. On Vicuna, the attacks converge to $100$ $ASR$ very quickly, noticing little to no difference between our initialization and the standard one.
Refer to caption Refer to caption Refer to caption
Figure 11: Comparison of $K\text{-}CRI$ ($K=1,5,25$) to standard and random initialization on the $GCG$ attack over the $AdvBench$ dataset. The attacks’ $ASR$ are presented on Llama-2 (left), Vicuna (middle), and Llama-3 (right).
Refer to caption Refer to caption Refer to caption
Figure 12: Comparison of $25\text{-}CRI$ to standard initialization on the $AutoDAN\text{-}GA$ attack over the $AdvBench$ dataset. The attacks’ $ASR$ are presented on Llama-2 (left), Vicuna (middle), and Llama-3 (right).
Refer to caption Refer to caption Refer to caption
Figure 13: Comparison of $25\text{-}CRI$ to standard initialization on the $AutoDAN\text{-}HGA$ attack over the $AdvBench$ dataset. The attacks’ $ASR$ are presented on Llama-2 (left), Vicuna (middle), and Llama-3 (right).
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 14: Embedding attack results on $AdvBench$. We compare $K\text{-}CRI$ ($K{=}1,5,25$) to standard and random initialization across five models. Top row (left to right): Llama-2, Vicuna, Llama-3. Bottom row (left to right): Llama-3.1, Mistral-7B.
C.4.4 LFS VS. Steps to Success - HarmBench
Figure 15: $LFS$ versus steps-to-success for four initialization strategies on Vicuna in the $GCG$ attack. The dashed line is the least-squares regression fit across all points (Pearson $r=0.31$, $p=1\times 10^{-5}$).
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 16: $LFS$ versus the number of optimization steps required for the embedding attack across five LLMs. Llama-2 ($r=0.62,\;p=4\times 10^{-17}$); Vicuna ($r=0.70,\;p=2\times 10^{-23}$); Llama-3-8B ($r=0.32,\;p=7\times 10^{-5}$); Llama-3.1-8B ($r=0.39,\;p=10^{-6}$); Mistral-7B ($r=0.51,\;p=4\times 10^{-11}$).
C.5 Ablation Study
C.5.1 Batch Size
In this part, we experiment with the batch size hyperparameter of the $GCG$ attack, which governs computational resources per attack iteration. A larger batch size allows considering more potential token replacements at each iteration. In Figure 17, we evaluate on the $HarmBench$ dataset, and in Figure 18, on $AdvBench$. We observe that for both datasets, using a small batch size of $16$, the $1\text{-}CRI$ initialization underperforms compared to other $CRI$s and the baseline. However, performance significantly improves, becoming comparable or superior to other initializations, when the batch size increases. Additionally, the baseline initialization severely underperforms relative to $25\text{-}CRI$ and $5\text{-}CRI$ at small batch sizes ($16$ and $32$), and only becomes somewhat comparable, yet still inferior, at larger batch sizes ($64$). This highlights the significance of our framework under varying computational resources.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 17: Comparison of $K\text{-}CRI$ ($K=1,5,25$) to standard and random initialization on the $GCG$ attack over the $HarmBench$ dataset. The attacks’ $ASR$ are presented on Llama-2 (left) and Vicuna (right). Across different batch sizes: 16 (top), 32 (center), and 64 (bottom).
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 18: Comparison of $K\text{-}CRI$ ($K=1,5,25$) to standard and random initialization on the $GCG$ attack over the $AdvBench$ dataset. The attacks’ $ASR$ are presented on Llama-2 (left) and Vicuna (right). Across different batch sizes: 16 (top), 32 (center), and 64 (bottom).
C.5.2 CRI Cost
Given a fine-tuning set $S_{FT}$ of size $N$, the total $CRI$ training cost is $C_{CRI}=\sum_{i=1}^{K}T_{i}$, where $T_{i}$ is the number of optimization steps for the $i$-th initialization in the set $\mathcal{T}_{K\text{-}CRI}$. At deployment, each test input uses this initialization to reduce optimization cost from $C_{\text{base}}$ (baseline) to $C_{\text{CRI-deploy}}$. The total amortized cost over $N_{test}$ test prompts is:
$C_{\text{total}}=C_{CRI}+N_{test}\cdot C_{CRI\text{-deploy}}\ll N_{test}\cdot C_{\text{base}}\quad\text{when }N_{test}\gg K.$
Ablation study of the number of steps used to create the initialization on $GCG$ in the $HarmBench$ dataset can be found in Figures 19, 20 and 21. We notice that quite consistently, the more steps used for training the $CRI$ set, the better the results in deployment are, with not a large difference between $400$ and $500$ initialization steps.
Refer to caption Refer to caption
Figure 19: Ablation on the training steps of the $25\text{-}CRI$ initialization, and the attack performance when used in deployment.
Refer to caption Refer to caption
Figure 20: Ablation on the training steps of the $5\text{-}CRI$ initialization, and the attack performance when used in deployment.
Refer to caption Refer to caption
Figure 21: Ablation on the training steps of the $1\text{-}CRI$ initialization, and the attack performance when used in deployment.
Cross-Dataset CRI Transferability
To evaluate the generalization capacity of $CRI$, we train the initialization set $\mathcal{T}_{K\text{-}CRI}$ on $S_{FT}^{\text{AdvBench}}$ and directly deploy it on $S_{FT}^{\text{HarmBench}}$ without retraining. Experiments on both Llama-2 and Vicuna show that $CRI$ maintains high $ASR$ across datasets, suggesting that the compliance-refusal structure captured by $CRI$ is model- and dataset-agnostic. This transferability further reduces the need for repeated initialization training, offering near-zero overhead in cross-dataset settings. In Figure 22, we compare using the same dataset for initialization and testing vs. using separate datasets. Here we test on the $HarmBench$ dataset, and compare using it for initialization vs. using $AdvBench$ for initialization.
Refer to caption Refer to caption
Figure 22: Attack success rates on $HarmBench$: in-dataset initialization versus cross-dataset initialization from $AdvBench$. Llama-2 (left) and Vicuna (right).
ASR Under Equalized Computational Budget
To better evaluate $CRI$’s effectiveness under fair cost conditions, we compare it to baseline attacks executed with a proportionally increased computational budget. Specifically, we benchmark $CRI$ against baselines that are allowed twice the number of optimization iterations (e.g., 500 steps). In Figure 23, we compared $25-CRI$ with baselines over a proportionally increased computational budget.
Figure 23: Comparison of $25-CRI$ and running the standard attack for an extended amount of steps.
. Results show that even under this relaxed constraint, $CRI$ consistently achieves higher attack success rates (ASR), demonstrating its efficiency in navigating the compliance-refusal space. This highlights $CRI$’s advantage not only in low-cost regimes, but also indeed enhances the $ASR$ performance of the attacks.
C.5.3 Additional Studies, Attacks Transfer, Defenses, Evaluations And Integration with other improvements
Initialization Transferability Between Attacks
We evaluated the transferability of initialization by extracting the 1-CRI initialization set from the GCG individual attack Liu et al. (2022) and using it to initialize the Embedding attack Schwinn et al. (2024), we choose the model $deepseek-llm-7b-chat$ as we find it new and robust for jailbreak attacks. While partial improvements were observed, the effectiveness was limited, suggesting that the optimization dynamics of these two distinct attack paradigms differ significantly. This highlights that while CRI can generalize across certain frameworks, its performance is sensitive to the underlying optimization mechanisms
Setup MSS ($\downarrow$) ASS ($\downarrow$) ASR (%) ($\uparrow$) LFS ($\downarrow$)
Baseline 20.0 20.04 100.0 2.2598
$1\text{-}CRI$ 10.0 10.20 100.0 1.6445
$25\text{-}CRI$ 13.0 12.96 100.0 1.3027
Transfer 14.5 14.32 100.0 2.2246
Table 7: Performance comparison of initialization strategies for deepseek-llm-7b-chat. Bold indicates best performance per metric.
Performance consistency: Integrating CRI with Advanced-Attack-Evaluations, integration with other Improvements-Methods and Defenses
Advanced Evaluation
In our work, we adhere to the evaluation setting used in the original attack baseline to ensure coherent results, however, we acknowledge that more sophisticated metrics could further enhance robustness. We have extended our evaluation as recommended to include a GPT-4-based judgment setting Liu et al. (2024a). Following the protocol from Liu et al. (2024a), GPT-4 is used to assess the $ASR$ of generated jailbreak prompts. Our results show that $CRI$ continues to improve $ASR$ in this setup. Furthermore, the best-performing initialization remains consistent across both the keyword-matching and GPT-4-based evaluation frameworks, reinforcing the reliability and transferability of $CRI$-enhanced prompts, the results are demonstrated in Figure 24.
Figure 24: Evaluation using GPT-4 as a judge.
BOOST
We present additional evaluations of our approach in more challenging settings. Specifically, we integrate BOOST Yu et al. into our attack framework, using a pre-trained initialization set in place of the standard $GCG$ initialization. The resulting variant—combining both $CRI$ and BOOST—achieves the highest $ASR$, demonstrating that $CRI$ is compatible with BOOST without requiring retraining. To compare against baselines in the setting proposed by the original BOOST authors, we use a keyword-matching evaluation method. In Figure 25, we compare our initializations to using BOOST Yu et al. .
Figure 25: Comparison of using our initialization, vs. using BOOST vs. using both BOOST and our initialization.
Perplexity Defense
When utilizing the same perplexity defense as in Liu et al. (2023).
We extended our evaluations to consider adversarial defenses. We present results on the perplexity-based filter suggested by Liu et al. (2023) on Llama-2, Vicuna and Llama-3. The perplexity-based defense does not affect the previously reported $ASR$ for the $AutoDAN$ attack, with or without $CRI$.
Warning: The following sections contains potentially offensive and harmful text.
Appendix D Results examples AutoDan-HGA
Figure 26: Examples of model generation by $AutoDAN\text{-}HGA$ on Llama-2 with and without $CRI$.
Figure 27: Examples of model generation by $AutoDAN\text{-}HGA$ on Llama-3 with and without $CRI$.
e
Appendix E Results examples GCG-M
Figure 28: Llama-2 responses to $GCG\text{-}M$ attacks.
e
Appendix F CRI Set Example
Figure 29: $CRI$ set for Llama-2 on the $GCG$ attack.
e